Nowlin pipe & steel has projected sales of 32,400 pipes this year, an ordering cost of $3 per order, and carrying costs of $1.50 per pipe.

A. What is the economic ordering quantity?

B. How many orders will be placed during the year?

C. What will be the average inventory?
Answer #1

Answer 1

EOQ =    2 * Demandof Pipes * CostperOrder Holding Cost

=   /2 * 32400 * 3 V 1.5

= 360

Answer 2

No of Order = Total Demand / EOQ

= 32400/360

= 90

Answer 3

Avg Inventory = EOQ/2

= 360/2 = 180

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.